# Multiple Stochastic Prompt Tuning for Practical Cross-Domain Few Shot Learning

### v1

@@ Abstract
In this work, we propose a practical cross-domain few-shot learning (pCDFSL) task, where a large-scale pre-trained model like CLIP can be easily deployed on a target dataset. The goal is to simultaneously classify all unseen classes under extreme domain shifts, by utilizing only a few labeled samples per class. The pCDFSL paradigm is source-free and moves beyond artificially created episodic training and testing regimes followed by existing CDFSL frameworks, making it more challenging and relevant to real-world applications. Towards that goal, we propose a novel framework, termed MIST (MultIple STochastic Prompt tuning), where multiple stochastic prompts are utilized to handle significant domain and semantic shifts. Specifically, multiple prompts are learnt for each class, effectively capturing multiple peaks in the input data. Furthermore, instead of representing the weights of the multiple prompts as point-estimates, we model them as learnable Gaussian distributions with two different strategies, encouraging an efficient exploration of the prompt parameter space, which mitigate overfitting due to the few labeled training samples. Extensive experiments and comparison with the state-of-the-art methods on four CDFSL benchmarks adapted to this setting, show the effectiveness of the proposed framework.
@@ 1 Introduction
Cross-domain few-shot learning (CDFSL) aims to adapt a deep learning model to novel classes with only a few labelled examples per class, where the target data has a significant distribution shift from the source data. Existing CDFSL approaches in literature [12, 9, 34] divide the source and target domain datasets into several episodes with fixed number of classes and training examples from each class (commonly known as N-way k-shot setting), to match training and testing conditions. These methods usually involve a resource-intensive meta-training on the source domain episodes followed by finetuning on the target domain episodes. Recent emergence of off-the-shelf large pretrained models like CLIP [24] offers a source-free alternative to replace the complex pretraining stage. Efficient transfer learning (ETL) techniques like prompt tuning [32, 15] have attempted such a setting focusing mainly on datasets with natural images and semantics, overlooking extreme domain shifts or specialized categories. Recently, few works have utilized CLIP for CDFSL [28, 4] for the episodic setup, where the model is fine-tuned on the episodes of the target dataset. However, in real-world, models are usually deployed where all the classes are present together, and rarely on isolated episodes with a few selected classes at a time.
Here, we propose a more realistic setting, practical cross-domain few-shot learning (pCDFSL), inspired by recent work on practical FSL [8], but for the more challenging CDFSL scenario, which unifies the CLIP-based few-shot ETL with the traditional CDFSL paradigms. In pCDFSL, a large-scale pre-trained model after being appropriately tuned using only few labeled samples from all the classes, can directly be deployed to the target dataset with extreme domain shifts. This setting better aligns with real-world deployment needs by enabling simultaneous classification across diverse, unseen classes without the episodic constraints. However, it poses significant challenges, as it involves learning all classes together with few labeled training samples, besides handling complex interactions among classes, leading to harder-to-discriminate class boundaries.
In this work, we propose a novel prompt learning framework for CLIP, termed MIST (MultIple STochastic Prompt tuning) to address these challenges. Analyzing the existing prompt tuning approaches, we observe that more often than not, unimodal modeling of class-wise data distributions does not hold. In contrast, our multiple prompts are better suited to capture the multimodal class distributions. Moreover, the availability of limited training samples also introduces the risk of overfitting. To this end, unlike standard optimization of prompt weights, here we model the different prompt parameters with different Gaussian distributions represented by learnable mean and variance vectors. This approach facilitates efficient exploration of the prompt parameter space, and prevents overfitting on the few training samples. The key contributions of this work can be summarized as follows:
@@
    We advocate the practical cross-domain few-shot learning (pCDFSL) task, which is more realistic and challenging compared to the constrained episodic setting.
@@
    We extensively evaluate the recent state-of-the-art CLIP-based methods for this task to generate strong baselines.
@@
    The novel MIST framework introduces multiple prompt vectors to effectively capture the multimodal distributions of the different classes.
@@
    We propose modeling the weights of the multiple prompts as learnable Gaussian distributions to mitigate overfitting on few samples and improve generalization.
@@
    Extensive experiments on standard CDFSL benchmarks adapted to this setting, demonstrate the effectiveness of the proposed MIST framework compared to state-of-the-art methods.
2 Related Works
Here, we briefly discuss the related work in literature. Cross-domain few-shot learning. The goal of CDFSL is to learn novel categories with a few training examples under large domain shifts. Existing methods in literature can be divided into two categories: meta-learning and transfer learning. Meta-learning approaches [12, 9, 35, 34], typically train a model to learn robust representations over a large number of source domain episodes, and then transfers this model to the target domain episodes. In [12], a simple pipeline with pretrained ViT is introduced, while [9] meta-trains a large-scale pretrained ViT by perturbing the original visual styles. Observing that CLS token of ViTs can absorb domain information, [35] proposes to decouple this information during source-domain training, before adapting on the target domain. [34] implements a new normalization layer to flatten loss landscapes in the representation space. Transfer learning methods [10, 19, 18] propose pretraining the model on source data followed by finetuning it on the target domain episodes. Recently, a few methods [28, 4] have explored adapting vision-language models like CLIP for CDFSL, by direct finetuning on the target episodes. In particular, [28] aims to enhance the robustness of CLIP by mixing domain-invariant text features with image features, while [4] introduces a regularization technique to increase the inter-class margin for enhanced class discrimination. However, all these approaches rely on the episodic setting, which involves a cumbersome training protocol, making them resource-intensive and impractical for real-world deployment.
@@ 2 Related Works
Prompt tuning of VLMs. Recently, vision-language models (VLMs) like CLIP [24] and ALIGN [13], which are pretrained on web-scale data have changed the landscape of computer vision with powerful zero-shot performance. However, adapting such large-scale models to downstream tasks is challenging, due to the risk of overfitting. Efficient transfer learning methods like prompt tuning [32, 15, 16] address this issue by optimizing only a few parameters added to these models. Specifically, CoOp [32] trains a few prompt vectors appended to the classname text, keeping the CLIP encoders frozen. MaPLe [15] proposes training prompts in both the textual and visual branches to improve multimodal alignment, while PromptSRC [16] further enhances performance by distilling knowledge from the frozen CLIP model. [29] incorporates class description features into the text encoder during prompt tuning to improve discriminability of the classifier. [20] tunes text prompts using a pool of diverse prompts for each class, while [7] adds learnable gaussian noise to each token of the learnable text prompt, to improve generalization. Our method utilizes a multiple prompt learning framework, whose weights are sampled from different learnable Gaussian distributions to address the pCDFSL task.
@@
Stochastic neural networks. Standard neural networks train weights deterministically as point-estimates. Contrarily, Bayesian Neural Networks [23, 3] model the weights as probability distributions, making them useful in handling uncertainty in predictions as well as learning robust representations. Stochastic classifiers have been explored in UDA [21], person re-identification [30], incremental learning [14] and DG [33] in prior literature. To the best of our knowledge, this is the first work which explores stochastic classifiers in the context of CDFSL.
3 Problem definition and Background
3.1 Practical CDFSL (pCDFSL) paradigm
Traditional CDFSL focuses on classifying novel categories using few samples under distribution shifts in an episodic manner. Let the source and target datasets be denoted as $D_{src}$ and $D_{tgt}$ respectively, which are split into a large number of tasks/ episodes $\mathcal{T}=(\mathcal{S},\mathcal{Q})$. The support set $\mathcal{S}=\{(X_{i},y_{i})\}^{N\times k}_{i=1}$ contains k samples from N (typically $5$) randomly chosen classes from all the classes in the dataset. The query set $\mathcal{Q}=\{(X_{i},y_{i})\}^{N\times q}_{i=1}$ contains q samples from the same set of N classes. This setup is commonly termed as the N-way k-shot setting.
In conventional CDFSL, the model is trained on $\mathcal{S}$ followed by evaluation on $\mathcal{Q}$ for each episode of the source dataset $D_{src}$. The methods are then evaluated based on the average accuracy and the variance over all the episodes of $D_{tgt}$. Here, we propose a more realistic practical CDFSL (pCDFSL) task, which is characterized as follows:
1) Source-free setup: A large pre-trained model like CLIP is utilized off-the-shelf, to efficiently leverage the pretrained knowledge base, without any access to source data.
2) Direct deployment: The model is adapted to the target dataset $D_{tgt}$ using k samples from all the classes simultaneously, instead of only the classes in the current episode. For example, if $D_{tgt}$ consists of $C$ categories, the model is trained on $k\times C$ datapoints.
3) Realistic evaluation: The fine-tuned model is tested on the entire test set of $D_{tgt}$. This is in contrast to the few query samples in each episode (typically $15$) in standard CDFSL, which does not reflect the difficulty that may be encountered while deploying models in the wild.
4) More informative results: In this setup, the model performance is not dependent on episodic difficulty. In standard CDFSL, some episodes can have less overlap among the class concepts and are easily separable, while others can have semantically similar classes which are difficult to discriminate, often leading to high variance across episodes.
@@ 3.2 Preliminaries
Base Network of MIST: In MIST, we employ a multimodal prompt learning strategy, where learnable prompt vectors are appended to the image and textual input branches. Specifically, let the learnable text prompt vectors be denoted as $\theta_{t}=\{\theta_{t_{1}},\theta_{t_{2}},...,\theta_{t_{m}}\}$ and the visual prompts as $\theta_{v}=\{\theta_{v_{1}},\theta_{v_{2}},...,\theta_{v_{m}}\}$. These are appended to the input text and image patches to form the modified inputs as: $\tilde{X}_{t}=\{t_{SOS},\theta_{t_{1}},...,\theta_{t_{m}},t_{1},t_{2},...,t_{CLS},t_{EOS}\}$ and $\tilde{X}_{v}=\{e_{CLS},\theta_{v_{1}},...,\theta_{v_{m}},e_{1},e_{2},...,e_{M}\}$ respectively. The extracted feature embeddings from the CLIP encoders are now $\tilde{z}_{t}=\mathcal{F}_{t}(\tilde{X}_{t})$ and $\tilde{z}_{v}=\mathcal{F}_{v}(\tilde{X}_{v})$. Here, the trainable textual prompts are passed through a learnable projection layer $f_{\phi}$ to obtain the visual prompts, i.e., $\theta_{v}=f_{\phi}(\theta_{t})$. Along with adding prompts to the inputs, we also adopt a deep prompting approach [15, 16], where learnable prompt vectors are attached after every transformer block. When adapting to a downstream task, these multimodal prompts are trained in an end-to-end manner, keeping the CLIP encoders frozen.
@@ 4 The Proposed Framework
In pCDFSL, the goal is to efficiently adapt the large-scale pre-trained CLIP using very few samples from all the classes simultaneously, under extreme domain and semantic shifts. Towards this goal, we propose two novel additions to the base network to handle different aspects of this setting: (i) Stochastic prompt learning to mitigate the risk of overfitting due to very few training examples and (ii) Multiple prompts to address the implicit unimodal assumption regarding the class data distributions. Together they constitute our final proposed framework, which is illustrated in Figure 3. We describe the motivation and details of the proposed additions next.
@@
Table 1: Stochastic prompt learning with two different sampling techniques on EuroSAT and ISIC. The fixed mean approach ($\mu,\sigma^{*}$) performs better in low shots, while sampling from a fully learnable distribution ($\mu^{*},\sigma^{*}$) performs better in higher shots.
@@ 4.1 Stochastic Prompt Learning
Although existing prompt tuning methods optimize a few learnable parameters appended to the inputs, the limited availability of training samples makes large-scale models like CLIP prone to overfitting as observed by [16]. While prior methods have addressed this using regularization modules [16] or architectural modifications [15], here we explore a different strategy using stochastic prompt learning to address the overfitting issue.
The main idea is to learn a distribution over the prompt parameters instead of optimizing them as point-estimates as is the standard practice. This helps to mitigate the uncertainty arising from scarce data, since each distinct sampled weight from this learnable distribution forms diverse decision boundaries for the few shot data, by allowing a richer exploration of the prompt parameter space. This provides an implicit regularization to the model without additional loss functions leading to more robust decision boundaries for the few-shot training samples. Specifically, we sample the text prompt weights from a Gaussian distribution $\mathcal{N}(\mu,\sigma)$, parameterized by mean $\mu$ and variance $\sigma$ as follows: $\theta_{t}\sim\mathcal{N}(\mu,\sigma)$. After every iteration, we backpropagate the loss to the learnable $\mu$ and $\sigma$ parameters. For end-to-end training, we use the Gaussian reparameterization trick [17] as follows:
@@
We illustrate the effect of stochastic prompt learning in Figure 1. In the deterministic case, passing the same examples through the trained model always projects them to the same points in the feature space. In contrast, the projections in the stochastic scenario spreads over a broader region, due to sampling of weights from the learned distribution. This variability implicitly encourages a margin between the class-specific features, resulting in more discriminative decision boundaries.
Figure 2: t-SNE visualization of image embeddings from MaPLe [15] (left) and TCP [29] (right) for the EuroSAT dataset. The classwise data distribution violates the unimodal assumption, due to high interclass similarity and intraclass variations.
To verify this phenomenon quantitatively, we consider two distinct strategies for sampling the text prompt parameters. In the first case, we fix the mean of the Gaussian to the standard prompt “A photo”, keeping the variance learnable. In the second case, both the mean and variance are learnable. The results on two representative datasets from the BSCDFSL [10] benchmark are shown in Table 1. We observe that in the low-shot setting, the first approach outperforms the second, while an opposite trend is seen in the higher-shot setting. This suggests that with very few training data (1-shot), directly optimizing the parameters of a distribution is challenging, but exploring variations around a standard prompt improves performance. On the contrary, learning the full distribution with more data outperforms the fixed mean approach. Hence, the two strategies complement each other in facilitating an efficient coverage of the prompt parameter space.

### v2

# Multiple Stochastic Prompt Tuning for Few-shot Adaptation under Extreme Domain Shift

@@ Abstract
Foundation Vision-Language Models (VLMs) like CLIP exhibit strong generalization capabilities due to large-scale pretraining on diverse image-text pairs. However, their performance often degrades when applied to target datasets with significant distribution shifts in both visual appearance and class semantics. Recent few-shot learning approaches adapt CLIP to downstream tasks using limited labeled data via adapter or prompt tuning, but are not specifically designed to handle such extreme domain shifts. Conversely, some works addressing cross-domain few-shot learning consider such domain-shifted scenarios but operate in an episodic setting with only a few classes per episode—limiting their applicability to real-world deployment, where all classes must be handled simultaneously. To address this gap, we propose a novel framework, MIST (Multiple Stochastic Prompt Tuning), for efficiently adapting CLIP to datasets with extreme distribution shifts using only a few labeled examples, in scenarios involving all classes at once. Specifically, we introduce multiple learnable prompts per class to effectively capture diverse modes in visual representations arising from distribution shifts. To further enhance generalization, these prompts are modeled as learnable Gaussian distributions, enabling efficient exploration of the prompt parameter space and reducing overfitting caused by limited supervision. Extensive experiments and comparisons with state-of-the-art methods demonstrate the effectiveness of the proposed framework.
@@ 1 Introduction
Foundation Vision-Language Models (VLMs) such as CLIP clip and ALIGN align have significantly advanced computer vision by enabling strong zero-shot generalization across tasks. Trained on large-scale image-text pairs, they learn robust representation spaces. However, such pretraining is costly and impractical to replicate for every new task, prompting growing interest in efficient adaptation using only a few labeled examples.
However, few-shot adaptation of such large-scale models is challenging due to the risk of overfitting and the potential loss of their original pretrained generalization capabilities. To address this, an emerging body of work has explored parameter-efficient fine-tuning techniques, such as prompt tuning coop and adapter tuning clip_adapter . These methods avoid full model fine-tuning by introducing a small set of trainable parameters while keeping the backbone frozen—either by modifying the input space (prompt tuning) or the output layers (adapter tuning). Despite these advances, such approaches primarily focus on standard benchmark datasets with natural images and class semantics (e.g., ImageNet, Caltech101), and largely overlook scenarios involving extreme domain shifts. In real-world applications, downstream datasets can exhibit significant shifts in both visual appearance and label semantics. For example, medical image datasets often feature domain-specific content and class names that do not align with natural image concepts and are typically unavailable for pretraining due to privacy concerns. While recent works have applied CLIP to cross-domain few-shot learning (CDFSL) semantic_guided , they typically adopt a source-free meta-testing setup, adapting to episodes with a few sampled classes (e.g., 5-way) from the target domain. Performance is averaged over many such episodes. However, this approach is computationally expensive and misaligned with real-world settings, where all target classes are present simultaneously.
In this work, we propose a novel prompt learning framework, MIST (Multiple Stochastic Prompt Tuning), for adapting CLIP to a more realistic setting, where target datasets exhibit significant domain and semantic shifts, and only a few labeled examples from all classes are available simultaneously. We first observe that extreme distribution shifts can lead to fragmented visual representations, forming separate and inconsistent clusters in the embedding space (Fig. 2). Moreover, large number of classes with semantic shifts (different class labels) can cause multiple class features to cluster together, resulting in ambiguous decision boundaries. To address these challenges, we introduce multiple learnable prompts per class, enabling better modeling of multi-modal feature distributions. Further, instead of directly optimizing prompt weights, we represent each prompt as a Gaussian distribution with learnable mean and variance, promoting diverse and well-separated representations while mitigating overfitting through efficient exploration of the prompt space. The key contributions of this work are summarized below:
@@
    We propose a novel framework for few-shot adaptation of CLIP to realistic scenarios involving extreme domain and label semantic shifts, with all target classes present simultaneously.
@@
    We analyze limitations of existing prompt tuning methods for few-shot setting, under severe distribution shifts.
@@
    We propose MIST, a novel prompt tuning framework that uses multiple class-specific prompts to model multimodal visual feature distributions.
@@
    We further represent each prompt as a learnable Gaussian distribution, enabling better generalization and reducing overfitting in low-data regimes.
@@
    Extensive experiments demonstrate that MIST outperforms state-of-the-art methods across multiple challenging benchmarks.
2 Related Work
Here, we briefly discuss the related work in literature.
Vision-language foundation models. Recently, foundation Vision-Language Models (VLMs) clip ; align ; blip have shown strong zero-shot generalization by learning aligned visual-textual representations from web-scale image-text pairs. However, their performance can degrade on domain-specific tasks with distribution shifts or rare class semantics not seen during pretraining. As retraining such large models for new tasks is impractical, recent efforts focus on efficient adaptation using limited labeled samples from the target domain.
@@ 2 Related Work
Few-shot adaptation of VLMs. Adapting these large-scale models to downstream tasks with few labeled training data is often challenging, due to the risk of overfitting. Efficient transfer learning methods like prompt tuning coop ; maple ; promptsrc or adapter tuning clip_adapter ; tip_adapter address this issue by optimizing only a few parameters added to these models, either in the input space or in the intermediate or output layers. For instance, CLIP-Adapter clip_adapter trains a classifier on the visual output features to modify the visual feature space, and Tip-Adapter tip_adapter stores the few-shot image prototypes, which are used to compute similarity with the test samples to guide the visual encoder. In prompt-tuning, CoOp coop trains a few prompt vectors appended to the classname text, keeping the CLIP encoders frozen. MaPLe maple proposes training prompts in both the textual and visual branches to improve multimodal alignment, while PromptSRC promptsrc further enhances performance by distilling knowledge from the frozen CLIP model. tcp incorporates class description features into the text encoder during prompt tuning to improve discriminability of the classifier. proda tunes text prompts using a pool of diverse prompts for each class, while bayesian_pl adds learnable gaussian noise to each token of the learnable text prompt, to improve generalization. These works mainly focus on standard image datasets with natural images and class semantics, often overlooking more realistic deployment scenarios where the target datasets may exhibit substantial domain shifts or unfamiliar, specialized label semantics. Some recent works have employed CLIP for the cross-domain few-shot learning (CDFSL) setup, which incorporates these challenges semantic_guided ; prompt_free . However, these methods take the episodic paradigm, which samples fixed number of classes in each episode, which is often unrealistic in real-world deployment.
@@
Stochastic neural networks. Standard neural networks train weights deterministically as point-estimates. Contrarily, Bayesian Neural Networks bnns ; weight_uncertainty model the weights as probability distributions, making them useful in handling uncertainty in predictions as well as learning robust representations. Stochastic classifiers have been explored in UDA star , person re-identification re-id , incremental learning s3c and DG stylematch in prior literature. To the best of our knowledge, this is the first work which explores stochastic classifiers for few-shot adaptation of VLMs under extreme domain shifts.
3 Problem Formulation
Given few labeled training examples from a target dataset, the task is to adapt the CLIP model efficiently to this data. Formally, we have a support set $\mathcal{S}=\{(X_{i},y_{i})\}^{C\times k}_{i=1}$ from the target dataset containing $k$ samples from all the $C$ classes simultaneously. Here, $y_{i}\in\{0,1\}^{C}$ is the corresponding ground truth label, and $k=\{1,2,4,8,16\}$, denotes the number of shots. The evaluation is done on the full test set. Here, in addition to the few-shot problem, the target dataset contains significant domain and label semantic shift from natural image datasets.
@@ 3.1 Preliminaries
Base Network of MIST: In MIST, we employ a multimodal prompt learning strategy, where learnable prompt vectors are appended to the image and textual input branches. Specifically, let the learnable text prompt vectors be denoted as $\theta_{t}=\{\theta_{t_{1}},\theta_{t_{2}},...,\theta_{t_{m}}\}$ and the visual prompts as $\theta_{v}=\{\theta_{v_{1}},\theta_{v_{2}},...,\theta_{v_{m}}\}$. These are appended to the input text and image patches to form the modified inputs as: $\tilde{X}_{t}=\{t_{SOS},\theta_{t_{1}},...,\theta_{t_{m}},t_{1},t_{2},...,t_{CLS},t_{EOS}\}$ and $\tilde{X}_{v}=\{e_{CLS},\theta_{v_{1}},...,\theta_{v_{m}},e_{1},e_{2},...,e_{M}\}$ respectively. The extracted feature embeddings from the CLIP encoders are now $\tilde{z}_{t}=\mathcal{F}_{t}(\tilde{X}_{t})$ and $\tilde{z}_{v}=\mathcal{F}_{v}(\tilde{X}_{v})$. Here, the trainable textual prompts are passed through a learnable projection layer $f_{\phi}$ to obtain the visual prompts, i.e., $\theta_{v}=f_{\phi}(\theta_{t})$. Along with adding prompts to the inputs, we also adopt a deep prompting approach maple ; promptsrc , where learnable prompt vectors are attached after every transformer block. When adapting to a downstream task, these multimodal prompts are trained in an end-to-end manner, keeping the CLIP encoders frozen.
@@ 4 The Proposed Framework
The goal is to efficiently adapt the large-scale pre-trained CLIP using very few samples from all the classes simultaneously, under extreme domain and semantic shifts. To this end, we present MIST (Fig. 3), which augments the base network with two novel modules to tackle these challenges: (i) Stochastic prompt learning to mitigate the risk of overfitting due to very few training examples and (ii) Multiple prompts to address the implicit unimodal assumption regarding the class data distributions, which we descrie below.
@@
Table 1: Stochastic prompt learning with two different sampling techniques on EuroSAT and ISIC datasets. The fixed mean approach ($\mu,\sigma^{*}$) performs better in low shots, while sampling from a fully learnable distribution ($\mu^{*},\sigma^{*}$) performs better in higher shots.
Figure 2: t-SNE visualization of image embeddings from MaPLe maple (left) and TCP tcp (right) for the EuroSAT dataset. The classwise data distribution violates the unimodal assumption, due to high interclass similarity and intraclass variations.
@@ 4.1 Stochastic Prompt Learning
Although existing prompt tuning methods optimize a few learnable parameters appended to the inputs, the limited availability of training samples makes large-scale models like CLIP prone to overfitting as observed by promptsrc . Additionally, when all target classes are present during adaptation, domain-specific or unfamiliar classnames can cause the class-specific features to lie closer together in the representation space, making inter-class boundaries harder to distinguish. To address these issues, we explore a novel strategy using stochastic prompt learning, where we model the prompts with learnable distributions.
The main idea is to learn a distribution over the prompt parameters instead of optimizing them as point-estimates as is the standard practice. This helps to mitigate the uncertainty arising from scarce data, since each distinct sampled weight from this learnable distribution forms diverse decision boundaries for the few shot data, by allowing a richer exploration of the prompt parameter space. This provides an implicit regularization to the model without additional loss functions leading to more robust decision boundaries for the few-shot training samples. Specifically, we sample the text prompt weights from a Gaussian distribution $\mathcal{N}(\mu,\sigma)$, parameterized by mean $\mu$ and variance $\sigma$ as follows: $\theta_{t}\sim\mathcal{N}(\mu,\sigma)$. After every iteration, we backpropagate the loss to the learnable $\mu$ and $\sigma$ parameters. For end-to-end training, we use the Gaussian reparameterization trick var_auto_encoder as follows:
@@
We observe from Fig. 1 that in the deterministic case, passing the same examples through the trained model always projects them to the same points in the feature space. In contrast, the projections in the stochastic scenario spreads over a broader region, due to sampling of weights from the learned distribution. This variability implicitly encourages a margin between the class-specific features, resulting in more discriminative decision boundaries.
To verify this, we consider two distinct strategies for sampling the text prompt parameters. First, we fix the mean of the Gaussian to the standard prompt “A photo”, keeping the variance learnable. In the second case, both the mean and variance are learnable. The results on two datasets bscdfsl are shown in Table 1. We observe that in low-shot setting, the first approach outperforms the second, while an opposite trend is seen in the higher-shot setting. This suggests that with very few training data (1-shot), directly optimizing the parameters of a distribution is challenging, but exploring variations around a standard prompt improves performance. On the contrary, learning the full distribution with more data outperforms the fixed mean approach. Hence, the two strategies complement each other in facilitating an efficient coverage of the prompt parameter space.